OpenStudy (anonymous):

9x−31 <= f(x) <= x^2+3x−22 determine lim as x goes to 3 what theorem should i use?

OpenStudy (amistre64):

its between the 2 :)

OpenStudy (anonymous):

Squeeze Theorem ? limit will be 4 ?

OpenStudy (dumbcow):

well both sides give a limit of -4 at x=3 i think its the squeeze theorem

OpenStudy (anonymous):

-4 i mean

OpenStudy (anonymous):

yup, it's right! squeez with lim -4 ^^ ty

OpenStudy (amistre64):

27-31 = -4 9+9-22 = 18-22 = -4

OpenStudy (anonymous):

yup (: